In a class, 70% of all students are boys and the rest are girls.

• 10% of the girls are smokers

• 20% of the boys are smokers

1) Find the probability that a randomly selected student is a smoker.



1
Expert's answer
2022-05-19T19:02:44-0400

There are 30% girls in the class (because 70% of all students are boys).

  • 10 percent of 30 is "30*10\/100=3" % (girls smokers)
  • 20 percent of 70 is "70*20\/100=14" % (boys smokers)

So, the probability that a randomly selected student is a smoker is "3+14=17\%" "=0.17"